# DiffuCoder: Understanding and Improving Masked Diffusion Models for Code Generation

### v1

@@ Abstract
Diffusion large language models (dLLMs) are compelling alternatives to autoregressive (AR) models because their denoising models operate over the entire sequence. The global planning and iterative refinement features of dLLMs are particularly useful for code generation. However, current training and inference mechanisms for dLLMs in coding are still under-explored. To demystify the decoding behavior of dLLMs and unlock their potential for coding, we systematically investigate their denoising processes and reinforcement learning (RL) methods. We train a 7B dLLM, DiffuCoder, on 130B tokens of code. Using this model as a testbed, we analyze its decoding behavior, revealing how it differs from that of AR models: (1) dLLMs can decide how causal their generation should be without relying on semi-AR decoding, and (2) increasing the sampling temperature diversifies not only token choices but also their generation order. This diversity creates a rich search space for RL rollouts. For RL training, to reduce the variance of token log-likelihood estimates and maintain training efficiency, we propose coupled-GRPO, a novel sampling scheme that constructs complementary mask noise for completions used in training. In our experiments, coupled-GRPO significantly improves DiffuCoder’s performance on code generation benchmarks (+4.4% on EvalPlus) and reduces reliance on AR causal during decoding. Our work provides deeper insight into the machinery of dLLM generation and offers an effective, diffusion-native RL training framework. https://github.com/apple/ml-diffucoder
@@ 1 Introduction
To leverage the benefits of non-autoregressiveness in dLLMs, it is important to understand how non-autoregressive the behavior of current dLLMs actually is. To this end, we introduce local and global autoregressive-ness (AR-ness) metrics (§4.1) to measure how closely their generation follows a left-to-right pattern. Our analysis reveals that dLLMs exhibit an entropy sink phenomenon (§4.2), which causes a strong causal bias during conditional generation using low-confidence remask decoding (Chang et al., 2022). We show that DiffuCoder can automatically decide how non-autoregressive it needs to be during decoding. When the sampling temperature is increased from the default $0.2$ to $1.2$, DiffuCoder becomes more flexible in its token generation order, freeing itself from strict left-to-right constraints, as Figure 1(a) shows. Unlike AR models, which primarily diversify token choices at higher temperatures, dLLMs additionally diversify the position of the generated token. With this increased diversity, DiffuCoder achieves higher pass@10 accuracy by changing the sampling temperature from $0.2$ to $1.2$ in our experiments (§4.3). The gain in pass@10 indicates the potential capacity of DiffuCoder, suggesting it can benefit from effective RL training to “elicit out” the most successful rollout samples.
@@ 2.1 Mask Diffusion Models
In diffusion models (Ho et al., 2020; Song et al., 2021), the forward process $q(\bm{x}_{1:T}|\bm{x}_{0})=\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})$ gradually corrupts data $\bm{x}_{0}\sim p_{data}(\bm{x}_{0})$ into noisy variables $\bm{x}_{1:T}$. The backward process models the joint probability as $p_{\theta}(\bm{x}_{0:T})=p_{\theta}(\bm{x}_{T})\prod_{t=1}^{T}p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$, denoising $\bm{x}_{t}$ to reconstruct $\bm{x}_{0}$. Discrete diffusion models (Ho et al., 2020; Zheng et al., 2024) define the forward process with a categorical distribution $q(\bm{x}_{t}|\bm{x}_{t-1})=\text{Cat}(\bm{x}_{t};\bm{Q}_{t}^{\top}\bm{x}_{t-1})$, where $\bm{x}_{t}\in\{0,1\}^{K}$ is a one-hot vector with vocabulary size $K$, and $\bm{Q}_{t}\in[0,1]^{K\times K}$ is the transition matrix where $[\bm{Q}_{t}]_{ij}$ represents the probability of transition from state $i$ to $j$. For absorbing discrete diffusion (Austin et al., 2021a), $\bm{Q}_{t}=(1-\beta_{t})I+\beta_{t}\mathbf{1}\bm{m}^{\top}$, where $\mathbf{1}$ is an all-one vector of size $K$ and $\bm{m}$ is the one-hot encoding of a special [MASK] token. The parameters $\theta$ are learned by minimizing the negative log-likelihood of $\bm{x}_{0}$ through the evidence lower bound (ELBO). For continuous time modeling (Shi et al., 2024; Sahoo et al., 2024), the discrete timesteps $t=\{1\dots T\}$ are scaled to a mask ratio within $[0,1]$, yielding the final ELBO at a sampled time $t$ as a weighted cross-entropy loss:
@@ 5 Coupled-GRPO
Table 2, together with Figure 7, demonstrates the effectiveness of coupled-GRPO training. In contrast, the baseline variants: d1, full-mask completion, and decoupled sampling, exhibit unstable reward learning. The rollout sampling temperature is also critical: as shown in Figure 6, DiffuCoder-Instruct attains a higher pass@10 at temperature $1.2$ than at $1.0$, mirroring the trend observed during coupled-GRPO training. Notably, RL fine-tuning shifts the optimal sampling temperature from $0.2$ to a larger value, such as $0.3$ or $0.4$, during evaluation, indicating that training sharpens the per-token distribution. This finding aligns with recent results for AR LLMs (Cui et al., 2025; Liu et al., 2025a; Agarwal et al., 2025), suggesting that the approaches proposed in these works may also be generalizable to dLLMs. Finally, at the new optimal temperature, global decoding AR-ness decreases, as shown in Figure 4 (right). A more interesting finding is that, as shown in Figure 1(c), when we use 0.5×0.5$\times$0.5 × fewer decoding steps (equivalent to a 2×2$\times$2 × generation speedup), training with coupled-GRPO results in a smaller performance drop compared to the model before training, suggesting that AR-ness is reduced and pallelism is increased. Detailed discussions are provided in Appx. 11 and Appx. D.
@@ 6 Related Work
Code generation is a crucial domain for LLMs (Roziere et al., 2023), exemplified by state-of-the-art open-source models like Qwen-2.5-Coder (Hui et al., 2024) and OpenCoder (Huang et al., 2024), with wide applications in areas such as coding assistants and agents (Xu et al., 2024). CodeFusion (Singh et al., 2023) was the first to combine diffusion models with code generation, but it was limited to small-scale models and simple tasks. Recent commercial-scale dLLMs, such as Mercury (Inception Labs et al., 2025) and Gemini (DeepMind, 2025), have demonstrated that diffusion-based code generators can achieve performance comparable to leading autoregressive code models while offering significantly faster generation speeds.

### v2

@@ Abstract
Diffusion large language models (dLLMs) are compelling alternatives to autoregressive (AR) models because their denoising models operate over the entire sequence. The global planning and iterative refinement features of dLLMs are particularly useful for code generation. However, current training and inference mechanisms for dLLMs in coding are still under-explored. To demystify the decoding behavior of dLLMs and unlock their potential for coding, we systematically investigate their denoising processes and reinforcement learning (RL) methods. We train a 7B dLLM, DiffuCoder, on 130B tokens of code. Using this model as a testbed, we analyze its decoding behavior, revealing how it differs from that of AR models: (1) dLLMs can decide how causal their generation should be without relying on semi-AR decoding, and (2) increasing the sampling temperature diversifies not only token choices but also their generation order. This diversity creates a rich search space for RL rollouts. For RL training, to reduce the variance of token log-likelihood estimates and maintain training efficiency, we propose coupled-GRPO, a novel sampling scheme that constructs complementary mask noise for completions used in training. In our experiments, coupled-GRPO significantly improves DiffuCoder’s performance on code generation benchmarks (+4.4% on EvalPlus) and reduces reliance on AR bias during decoding. Our work provides deeper insight into the machinery of dLLM generation and offers an effective, diffusion-native RL training framework. https://github.com/apple/ml-diffucoder
@@ 1 Introduction
To leverage the benefits of non-autoregressiveness in dLLMs, it is important to understand how non-autoregressive the behavior of current dLLMs actually is. To this end, we introduce local and global autoregressive-ness (AR-ness) metrics (§4.1) to measure how closely their generation follows a left-to-right pattern. Our analysis reveals that dLLMs exhibit an entropy sink phenomenon (§4.2), which causes a strong causal bias during conditional generation using low-confidence remasking decoding (Chang et al., 2022). We show that DiffuCoder can automatically decide how non-autoregressive it needs to be during decoding. When the sampling temperature is increased from the default $0.2$ to $1.2$, DiffuCoder becomes more flexible in its token generation order, freeing itself from strict left-to-right constraints, as Figure 1(a) shows. Unlike AR models, which primarily diversify token choices at higher temperatures, dLLMs additionally diversify the position of the generated token. With this increased diversity, DiffuCoder achieves higher pass@10 accuracy by changing the sampling temperature from $0.2$ to $1.2$ in our experiments (§4.3). The gain in pass@10 indicates the potential capacity of DiffuCoder, suggesting it can benefit from effective RL training to “elicit out” the most successful rollout samples.
@@ 2.1 Mask Diffusion Models
In diffusion models (Ho et al., 2020; Song et al., 2021), the forward process $q(\bm{x}_{1:T}|\bm{x}_{0})=\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})$ gradually corrupts data $\bm{x}_{0}\sim p_{data}(\bm{x}_{0})$ into noisy variables $\bm{x}_{1:T}$. The backward process models the joint probability as $p_{\theta}(\bm{x}_{0:T})=p_{\theta}(\bm{x}_{T})\prod_{t=1}^{T}p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$, denoising $\bm{x}_{t}$ to reconstruct $\bm{x}_{0}$. Discrete diffusion models (Hoogeboom et al., 2021; Zheng et al., 2024) define the forward process with a categorical distribution $q(\bm{x}_{t}|\bm{x}_{t-1})=\text{Cat}(\bm{x}_{t};\bm{Q}_{t}^{\top}\bm{x}_{t-1})$, where $\bm{x}_{t}\in\{0,1\}^{K}$ is a one-hot vector with vocabulary size $K$, and $\bm{Q}_{t}\in[0,1]^{K\times K}$ is the transition matrix where $[\bm{Q}_{t}]_{ij}$ represents the probability of transition from state $i$ to $j$. For absorbing discrete diffusion (Austin et al., 2021a), $\bm{Q}_{t}=(1-\beta_{t})I+\beta_{t}\mathbf{1}\bm{m}^{\top}$, where $\mathbf{1}$ is an all-one vector of size $K$ and $\bm{m}$ is the one-hot encoding of a special [MASK] token. The parameters $\theta$ are learned by minimizing the negative log-likelihood of $\bm{x}_{0}$ through the evidence lower bound (ELBO). For continuous time modeling (Shi et al., 2024; Sahoo et al., 2024), the discrete timesteps $t=\{1\dots T\}$ are scaled to a mask ratio within $[0,1]$, yielding the final ELBO at a sampled time $t$ as a weighted cross-entropy loss:
@@ 5 Coupled-GRPO
Table 2, together with Figure 7, demonstrates the effectiveness of coupled-GRPO training. In contrast, the baseline variants: d1, full-mask completion, and decoupled sampling, exhibit unstable reward learning. The rollout sampling temperature is also critical: as shown in Figure 6, DiffuCoder-Instruct attains a higher pass@10 at temperature $1.2$ than at $1.0$, mirroring the trend observed during coupled-GRPO training. Notably, RL fine-tuning shifts the optimal sampling temperature from $0.2$ to a larger value, such as $0.3$ or $0.4$, during evaluation, indicating that training sharpens the per-token distribution. This finding aligns with recent results for AR LLMs (Cui et al., 2025; Liu et al., 2025a; Agarwal et al., 2025), suggesting that the approaches proposed in these works may also be generalizable to dLLMs. Finally, at the new optimal temperature, global decoding AR-ness decreases, as shown in Figure 4 (right). A more interesting finding is that, as shown in Figure 1(c), when we use 0.5×0.5$\times$0.5 × fewer decoding steps (equivalent to a 2×2$\times$2 × generation speedup), training with coupled-GRPO results in a smaller performance drop compared to the model before training, suggesting that AR-ness is reduced and parallelism is increased (Wu et al., 2025). Detailed discussions are provided in Appx. 11 and Appx. D.
@@ 6 Related Work
Code generation is a crucial domain for LLMs (Roziere et al., 2023; Sun et al., 2024), exemplified by state-of-the-art open-source models like Qwen-2.5-Coder (Hui et al., 2024) and OpenCoder (Huang et al., 2024), with wide applications in areas such as coding assistants and agents (Xu et al., 2024). CodeFusion (Singh et al., 2023) was the first to combine diffusion models with code generation, but it was limited to small-scale models and simple tasks. Recent commercial-scale dLLMs, such as Mercury (Inception Labs et al., 2025) and Gemini (DeepMind, 2025), have demonstrated that diffusion-based code generators can achieve performance comparable to leading autoregressive code models while offering significantly faster generation speeds.